# Reward Poisoning Attack Against Offline Reinforcement Learning

### v1

@@ 3.1 Adversarial Reward Engineering
Under an attack within the adversarial reward engineering framework, any efficient learning algorithm satisfying Assumption 3.1 will learn the policy with a near-optimal performance on $\widehat{R}$: $J_{\widehat{R}}(\hat{\pi}^{*})\geq\max_{\pi\in\Pi_{\mu}}J_{\widehat{\mathcal{R}}}(\pi)-\delta$. The efficiency of the attack satisfies $V=J_{\mathcal{R}}(\hat{\pi}^{*})$, $B=\max_{i}|\widehat{\mathcal{R}}(s_{i},a_{i})-r_{i}|$ and $C=\sum_{i}|\widehat{\mathcal{R}}(s_{i},a_{i})-r_{i}|$
In general, the attacker needs to find a suitable $\widehat{\mathcal{R}}$ such that the required budget is limited, and the near-optimal policies in the adversarial reward function have poor performance in the true environment.
3.2 Inverted Reward Attack
Following the idea of reward engineering, one can immediately find an attack that can make the agent learn the worst possible policy under Assumption 3.1.
Definition 3.4 (Fully inverted reward attack).
The inverted reward attack flips the sign of all the rewards in the dataset. In Theorem 3.5 we show the efficiency of the fully inverted reward attack.
@@ 3.2 Inverted Reward Attack
Maximizing the cumulative reward on $-\mathcal{R}$ is equivalent to minimizing that on $\mathcal{R}$. Therefore, an efficient learning algorithm will learn the policy with the worst performance on the true reward function among the policies supported by the dataset. Despite achieving the worst outcome for the learner, the inverted reward attack faces the problem of requiring high budgets which is impractical. Since a practical attack should only corrupt a small portion of data, we consider an attack that can only randomly invert a part of the reward. We call this attack the ‘random inverted reward attack’. This attack strategy has been widely used in previous works as a non-trivial attack (Ye et al., 2023; Zhang et al., 2022).
Definition 3.6 (Random inverted reward attack).
3.3 Policy Contrast Attack
Here we introduce our attack. First, in Theorem 3.7 we give a sufficient condition and another necessary for the adversarial reward function to make the agent learn a low-performing policy. The two conditions are the same when $\delta$ becomes 00.
Theorem 3.7.
Consider an attack in the adversarial reward engineering framework with $-^\widehat{\mathcal{R}}$. To make the actual performance of the learned policy lower than a value $J_{\mathcal{R}}(\pi_{0})<V$, a sufficient condition is that the adversarial reward function satisfies
@@ 3.3 Policy Contrast Attack
In both cases, we call $\pi_{1}$ that satisfies the conditions ‘bad policy’ and $\pi_{2}$ ‘good policy’.
@@
Theorem 3.7 suggests that the attacker should make a low-performing policy have a higher performance, in the adversarial reward function, than any high-performing policies. Formally, denote $\Delta J(\pi)=J_{\widehat{\mathcal{R}}}(\pi)-J_{\mathcal{R}}(\pi)$. Then to make $J_{\widehat{\mathcal{R}}}(\pi_{1})>J_{\widehat{\mathcal{R}}}(\pi_{2})+\delta$, the attacker needs to ensure that $\Delta J(\pi_{1})-\Delta J(\pi_{2})>J(\pi_{2})-J(\pi_{1})+\delta$. In other words, the attacker needs to make $\Delta J(\pi_{1})$ high for some bad policy $\pi_{1}$ and $\Delta J(\pi_{2})$ low for any good policy $\pi_{2}$. Inspired by this, at a high level, the key idea for our attack is that in the adversarial reward function, some low-performing policies appear as high-performing (bad policies look good), while any high-performing policy appears as low-performing (good policies look bad). Based on the idea, we construct our attack called ‘policy contrast attack’.
Bad policies look good: Following the idea, the first step of the attacker is to find some low-performing policies supported by the dataset. To achieve this, the attacker can train with any efficient algorithm on the dataset with the inverted reward. By Theorem 3.5, the attacker can learn the worst-performing policy supported by the dataset, which is exactly the goal. Intuitively, this policy and the policies that behave similarly to it should have low performance in practice. Therefore, to make some low-performing policies have a relatively high performance in an adversarial environment, a straightforward way is to increase the reward associated with the actions close to that given by the worst-performing policies. For this purpose, we denote the distance measure between two actions $a_{1},a_{2}$ as $d_{a}(a_{1},a_{2})$. For the continuous action space as the ones considered in the continuous control settings (Fu et al., 2020), we choose the distance measure as $d_{a}(a_{1},a_{2})=||a_{1}-a_{2}||_{2}$.
Good policies look bad: The next step of the attacker is to find the high-performing policies supported by the dataset and make them have low performance in the adversarial reward function. A straightforward idea is to find a set of different good policies such that any good policy will behave similarly to one of them. Following this idea, we use an iterative learning approach to find good policies. The high-level idea is to learn a high-performing policy different from previously learned policies at each iteration. The algorithm to find good policies is described in Alg 1.
@@
In the iterative learning of good policies, the attacker starts with an empty set of good policies $\Pi_{2}=\emptyset$. At the first iteration, the attacker trains on the original dataset with an efficient learning algorithm and then adds the learned policy to the good policy set $\Pi_{2}$. In the subsequent iterations, the attacker modifies the dataset by decreasing the rewards at the actions close to one of the actions given by the policies from the good policy set $\Pi_{2}$. Note that this modification is not the corruption for the agent’s training. The attacker trains on the modified dataset and then adds the learned policy to the good policy set. Theorem 3.8 shows that policies of high performance will behave similarly to some policies from the good policy set.
Theorem 3.8.
Let $V$ be the performance of the learned good policy in the last iteration. For any policy $\pi\in\Pi_{\mu}$ such that $\min_{s\in\mathcal{S},\pi_{2}\in\Pi_{2}}d_{a}(\pi(s),\pi_{2}(s))>d$, its performance satisfies $J_{\mathcal{R}}(\pi)<V+\delta$.
To make the policies from the good policy set have bad performance in the adversarial reward, one can simply decrease the rewards associated with the actions close to the ones given by the good policies.
Algorithm 2 Policy Contrast Attack
  Params: distance threshold $d$, corruption parameters $\Delta_{1}$, $\Delta_{2}$, number of iterations $K$
  Initialize $\widehat{\mathcal{D}}=\{\}$.
  Learn a good policy set $\Pi_{2}$ through Alg 1 with $K$ iterations
  for $i=1$ to $N$ do
     Get state-action-reward $(s_{i},a_{i},r_{i})$ from $\mathcal{D}$
     if $\min_{\pi\in\Pi_{2}}d_{a}(a_{i},\pi(s_{i}))\leq d$ then
        Modify $\hat{r}_{i}=-\Delta_{2}$
     end if
     if $d_{a}(a_{i},\pi_{1}(s_{i})\leq d$ then
        Modify $\hat{r_{i}}=r_{i}+\Delta_{1}\cdot(1-d_{a}(a_{i},\pi_{1}(s_{i})/d)$
     end if
  end for
  Output: $\widehat{\mathcal{D}}$
Practical implementation: Finally, we consider the practical implementation of the attack method. First, the attacker finds a bad policy $\pi_{1}$ by training on the dataset with inverted rewards and a good policy set $\Pi_{2}$ by Alg 1. Intuitively, with a higher value of $K$ for the size of $\Pi_{2}$, more good policies of different behaviors will be covered. Empirically, we show that the performance of the attack is not sensitive to the value of $K$. Then, the attacker modifies the reward through an adversarial reward function that increases the rewards for bad actions and decreases the rewards for good actions. In Alg 2 we show the pseudo code for the policy contrast attack. The adversarial reward function in Alg 2 satisfies:
^$\widehat{\mathcal{R}}(s,a)=\begin{cases}\mathcal{R}(s,a)+\Delta_{1}\cdot\\ (1-\frac{d_{a}(a,\pi_{1}(s))}{d}),&\text{if }d_{a}(a,\pi_{1}(s_{2}))\leq d\\ -\Delta_{2},&\text{else if }\\ &\min_{\pi_{2}\in\Pi_{2}}d_{a}(a,\pi_{2}(s))\leq d\\ \mathcal{R}(s,a),&\text{otherwise}\end{cases}$
Here $\Delta_{1}$, $\Delta_{2}$, and $d$ are the hyper-parameters of the attack determined by the attack budgets. The first part of the reward function is to increase the rewards associated with the bad actions given by the bad policy. The formation of the reward modification in this part follows Xu & Singh (2023). The second part of the reward function is to decrease the rewards associated with good policies.
Note that there are some cases where the actions given by the good policies and the bad policy are close. In this case, there will be a conflict for the aforementioned corruption strategy. Intuitively, these cases are rare as usually good and bad actions should not be similar. In addition, in these cases, a change in reward has a similar influence on the performances of good and bad policies in the adversarial function, suggesting that the corruption in these cases has a limited impact on the ranking of good and bad policies in the adversarial reward. To ensure that the bad policy always has a higher performance in the adversarial environment, we always increase the reward associated with the bad actions whenever there is a conflict.
At last, we want to discuss why the policy contrast attack should be more efficient than the random inverted attack. At a high level, both the policy contrast and the inverted reward attack want to inverse the agent’s ranking of good and bad policies. In contrast to the inverted reward attack that treats all state actions equally and inverse the performance of all policies, the policy contrast attack focuses on policies that matter more as pointed out by Theorem 3.7 and the state actions that are more important. This potentially makes the policy contrast attack more efficient in terms of budget.
Figure 2: Influence of different $B$ budget on the attack.
Figure 3: Influence of different $C$ budget on the attack.
4 Experiments
4.1 Performance of Different Attack Strategy
We test the effectiveness of our attack against current state-of-the-art offline RL algorithms training on the standard D4RL offline RL benchmark’s continuous control domains (Fu et al., 2020). we show that with limited attack budgets, our policy contrast attack is in general more effective.
For the learning algorithms, we choose IQL  (Kostrikov et al., 2021), CQL (Kumar et al., 2020), and TD3-BC (Fujimoto & Gu, 2021), which are frequently used as baselines in current offline RL studies. The implementation of these algorithms is based on CORL library (Tarasov et al., 2022). Recall that the attacker also needs a learning algorithm to learn good and bad policies. To ensure that the attacker has no information about the learning agent, the learning agent and the attacker always use different learning algorithms.
For the baseline attack, we choose the random invert attack as it is the only non-trivial in the current literature. In the ablation study, we also show the result of using only one part of the attack strategy from the policy contrast attack, that is, the attack only makes good policy look bad or bad policy look good.
For the attack budget, following Ye et al. (2023), we always set the budget for $B$ and $C$ to be the amount required by the random inverted reward attack that corrupts $30\%$ of the data. We test the influence of different attack budgets on the efficiency of the attack in the ablation study.
For the hyper-parameters of the attack, the corruption parameters $\Delta_{1}$ and $\Delta_{2}$ are set to maximally utilize the budget $B$, and the attack radius $d$ is set to maximally utilize the budget $C$. The number of good policies, which is also the number of iterations $K$, is set as $5$ here. In the ablation study, we show that the efficiency of the algorithm is not sensitive to the choice of corruption parameters and the number of good policies. For each experiment, we run $5$ times with different random seeds and show the average result.
Results: The offline RL algorithms we choose iteratively update their learned policies. To intuitively show the details of their training process under the attacks, we evaluate and show the performance of the policy learned by the algorithm at each iteration during training. In the appendix, we show the quantified result of the learning algorithms’ performance under the attack.
Fig 1 shows the main results of running our policy contrast attack in different learning scenarios. The naive baseline is the training with no attack. The ideal effect of the attack is the training with fully inverted reward as the learning algorithm tries to learn the worst-performing policy in this case. We observe that our policy contrast attack is usually effective for any combination of datasets and learning algorithms while the random inverted attack has almost no effect on the learning result except for the Hopper-Medium Replay dataset where both attacks are similarly effective.

### v2

@@ 4.2 Adversarial Reward Engineering Framework
Let $\widehat{\mathcal{R}}$ be the adversarial reward of an instance of the adversarial reward engineering framework. Let $\hat{\Pi}^{*}:=\{\pi|\pi\in\Pi_{\mu},J_{\widehat{R}}(\hat{\pi}^{*})\geq\max_{\pi\in\Pi_{\mu}}J_{\widehat{\mathcal{R}}}(\pi)-\delta\}$ be the $\delta$-optimal supported policies in $\widehat{\mathcal{R}}$. The efficiency of the attack satisfies $V\leq\max_{\pi\in\hat{\pi}^{*}}J_{\mathcal{R}}(\pi)$, $B=\max_{i}|\widehat{\mathcal{R}}(s_{i},a_{i})-r_{i}|$ and $C=\sum_{i}|\widehat{\mathcal{R}}(s_{i},a_{i})-r_{i}|$.
The proof for all theorems and lemmas can be found in the Appendix. Recent works about robust offline RL widely consider an attack we call ‘random inverted reward attack’ for empirical evaluation against their learning algorithms Zhang et al. (2022); Ye et al. (2023); Li et al. (2024). Such an attack randomly flips the signs of a portion of rewards in the datasets and can be treated as an instance of the adversarial reward engineering framework. One can find a detailed analysis in the Appendix. Next, we show a way to construct a more efficient instance of the adversarial reward engineering framework and justify the design of our policy contrast attack.
4.3 Towards Efficient Adversarial Reward Engineering Attack: Policy Contrast Attack (PCA)
To ensure that the agent learns a low-performing policy, an efficient instance of the adversarial reward engineering attack needs to make the optimal or near-optimal supported policies in the adversarial environment have poor performance in the true environment. So, our goal is to find a way to construct such an adversarial reward model while limiting the required attack budget. However, it is still impractical to find all near-optimal supported policies and their performance in the true environment as the environment dynamics are unknown to the attacker. Therefore, we focus on finding a feasible way to construct a reward function such that any near-policy policies in the reward function are likely to perform poorly in the true environment.
In Theorem 4.5, we show a sufficient condition and a necessary condition for the adversarial reward function to make the agent learn a low-performing policy with performance less than $V$. The two conditions are the same when $\delta=0$.
@@ 4.3 Towards Efficient Adversarial Reward Engineering Attack: Policy Contrast Attack (PCA)
Consider an attack in the adversarial reward engineering framework with $\widehat{\mathcal{R}}$. To make the actual performance of the learned policy by an arbitrary $\delta$-optimal pessimistic learning algorithm lower than a value $J_{\mathcal{R}}(\pi_{0})<V$, a sufficient condition is that the adversarial reward function satisfies
@@
In both cases, we call $\pi_{1}$ that satisfies the conditions ‘bad policy’ and $\pi_{2}$ ‘good policy.’
@@
Theorem 4.5 suggests that the attacker should make a low-performing policy have a higher performance in the adversarial reward function than any high-performing policies. Formally, denote $\Delta J(\pi)=J_{\widehat{\mathcal{R}}}(\pi)-J_{\mathcal{R}}(\pi)$, to make $J_{\widehat{\mathcal{R}}}(\pi_{1})>J_{\widehat{\mathcal{R}}}(\pi_{2})+\delta$, the attacker needs to ensure that $\Delta J(\pi_{1})-\Delta J(\pi_{2})>J(\pi_{2})-J(\pi_{1})+\delta$. In other words, the attacker needs to make $\Delta J(\pi_{1})$ high for some bad policy $\pi_{1}$ and $\Delta J(\pi_{2})$ low for any good policy $\pi_{2}$. To limit the cost of the attack, the attacker should also determine the rewards to poison that have more influence on $\Delta J(\pi_{1})$ and $\Delta J(\pi_{2})$. This directly inspires the key idea for finding an efficient attack that in the adversarial reward function, the rewards for some low-performing policies are increased so that they appear as high-performing and vice versa.
In terms of finding the bad policies, we show in Lemma B.2 that the attacker can find nearly the worst policy supported by the dataset by training on the dataset of all rewards inverted.
Lemma 4.6.
When training on the dataset $\mathcal{D}$ with the signs of all rewards inverted, the learning agent will learn the policy supported by the dataset with nearly the worst performance $V\leq J_{\mathcal{R}}(\pi_{0})=\min_{\pi\in\Pi_{\mu}}J_{\mathcal{R}}(\pi)+\delta$.
In terms of finding the good policies, Lemma 4.7 shows that policies of high performance will behave similarly to some policies from the good policy set constructed through Alg 2.
Lemma 4.7.
Let $V$ be the performance of the learned good policy in the last iteration of Alg 2. For any policy $\pi\in\Pi_{\mu}$ such that $\min_{s\in\mathcal{S},\pi_{2}\in\Pi_{2}}d_{a}(\pi(s),\pi_{2}(s))>d$, its performance satisfies $J_{\mathcal{R}}(\pi)<V+\delta$.
After finding the good and bad policies, the final step is to make them look bad and good in the adversarial environment. This can be achieved straightforwardly by decreasing or increasing the rewards associated with the actions given by the good or bad policies.
Based on the discussion above, we formulate ‘policy contrast attack.’ The attack consists of two parts as follow:
  1. 1.
    Bad policies look good: The attacker trains on the training dataset with all rewards inverted and learns ‘bad’ policies. Then, the attacker poisons the dataset by increasing the rewards associated with the actions close to that given by the ‘bad’ policies.
  2. 2.
    Good policies look bad: The attacker iteratively learns a set of ‘good’ policies with Algorithm 2. Then, the attacker poisons the dataset by decreasing the rewards associated with the actions close to that given by the ‘good’ policies.
Formally, the policy contrast attack is given in Alg 1. Here $\Delta_{1}$, $\Delta_{2}$, and $d$ are the hyper-parameters of the attack and set to utilize given attack budgets $B$ and $C$ fully. $d_{a}$ is a measure of the distance between actions. For the actions from a continuous space, we choose $d_{a}(a_{1},a_{2})=||a_{1}-a_{2}||_{2}$. The corresponding adversarial reward function in Alg 1 satisfies:
^$\widehat{\mathcal{R}}(s,a)=\begin{cases}\mathcal{R}(s,a)+\Delta_{1}\cdot(1-\frac{d_{a}(a,\pi_{1}(s))}{d}),&\text{if }d_{a}(a,\pi_{1}(s_{2}))\leq d\\ -\Delta_{2},&\text{else if }\min_{\pi_{2}\in\Pi_{2}}d_{a}(a,\pi_{2}(s))\leq d\\ \mathcal{R}(s,a),&\text{otherwise}\end{cases}$
Algorithm 1 Policy Contrast Attack
Params: distance measure $d_{a}$, distance threshold $d$, corruption parameters $\Delta_{1}$, $\Delta_{2}$, number of iterations $K$
Initialize $\widehat{\mathcal{D}}=\{\}$.
Learn a good policy set $\Pi_{2}$ through Alg 2 with $K$ iterations
for $i=1$ to $N$ do
     Get state-action-reward $(s_{i},a_{i},r_{i})$ from $\mathcal{D}$
     if $\min_{\pi\in\Pi_{2}}d_{a}(a_{i},\pi(s_{i}))\leq d$ then
         Modify $\hat{r}_{i}=-\Delta_{2}$
     end if
     if $d_{a}(a_{i},\pi_{1}(s_{i})\leq d$ then
         Modify $\hat{r_{i}}=r_{i}+\Delta_{1}\cdot(1-d_{a}(a_{i},\pi_{1}(s_{i})/d)$
     end if
end for
Output: $\widehat{\mathcal{D}}$
The first part of the reward function is to increase the rewards associated with the bad actions given by the bad policy. The formation of the reward modification in this part follows Xu & Singh (2023). The penalty on the reward is a linear function depending on the distance between an action and the bad policy’s action. This is to make it easier for the agent to converge to the bad policy. The second part of the reward function is to decrease the rewards associated with good policies. The penalty here is simply making the reward a low value, as in this case, we merely want the agent to learn something different from the good policies.
Note that there are some cases where the actions given by the good and bad policies are close. There will be a conflict for the aforementioned corruption strategy in this case. Intuitively, these cases are less common as usually good and bad actions should not be similar. In addition, in these cases, a change in reward has a similar influence on the performances of good and bad policies in the adversarial function, suggesting that the corruption in these cases has a limited impact on the ranking of good and bad policies in the adversarial reward. To ensure that the bad policy always performs better in the adversarial environment, we always increase the reward associated with the bad actions whenever there is a conflict.
5 Experiments
5.1 Performance of Different Attack Strategy
We test the effectiveness of our attack against current state-of-the-art offline RL algorithms training on the standard D4RL offline RL benchmarkFu et al. (2020). We show that with limited attack budgets, our policy contrast attack is generally more effective.
For the learning algorithms, we choose IQL  Kostrikov et al. (2021), CQL Kumar et al. (2020), and TD3-BC Fujimoto & Gu (2021), which are frequently used as baselines in current offline RL studies. These algorithms are implemented based on CORL library Tarasov et al. (2022). Recall that the attacker also needs a learning algorithm to learn good and bad policies. To ensure that the attacker has no information about the learning agent, the learning agent and the attacker always use different learning algorithms.
We choose three baseline attacks. The first one is the random invert attack, the only non-trivial attack against offline RL in the current literature. The second and the third ones are the random policy invert (RPI) attack and random policy promote (RPP) attack proposed in Xu et al. (2022) for the online setting, as it is the only existing attack closely related to and can be adapted to work in the offline setting. In the appendix, we show the result of using only one part of the attack strategy from the policy contrast attack.
For the attack budget, following Ye et al. (2023), we always set the budget for $B$ and $C$ to be the amount required by the random inverted reward attack that corrupts $30\%$ of the data. In the ablation study, we tested the influence of different attack budgets on the attack’s efficiency.
For the hyper-parameters of the attack, the corruption parameters $\Delta_{1}$ and $\Delta_{2}$ are set to fully utilize the budget $B$, and the attack radius $d$ is set to fully use the budget $C$. The number of good policies, i.e., the number of iterations $K$, is set as $5$. In the appendix, we show that the algorithm’s efficiency is not sensitive to the choice of corruption parameters and the number of good policies. We run $5$ times for each experiment with different random seeds and show the average result.
Note that the construction of our attack does not depend on the assumption we made for the learning algorithms. Although the learning algorithms may not strictly satisfy the assumption of efficient learning algorithms, we show that our attack is still effective.
Main Results: The offline RL algorithms we choose iteratively update their learned policies. To straightforwardly show the effectiveness of attacks, we use the performance of the policy learned by the agent at the end of training to represent its learning efficiency. In the appendix, to intuitively show the details of the training process under the attacks, we show the performance of the policies learned by the algorithm at each iteration during training. We will also show the full training log for the experiments in the ablation studies.
D4RL Dataset Normal Fully inverted PCA Rand 30% RPP RPI
HalfCheetah-ME - TD3_BC 91.0 $\pm$ 2.9 13.1 $\pm$ 1.9 76.2 $\pm$ 2.1 89.8 $\pm$ 3.2 88.95 $\pm$ 2.65 81.3 $\pm$ 5.1
Hopper-ME - TD3_BC 105.9 $\pm$ 3.3 16.9 $\pm$ 6.3 82.1 $\pm$ 6.4 99.8 $\pm$ 8.9 88.22 $\pm$ 8.37 97.4 $\pm$ 6.9
HalfCheetah-MR - IQL 44.3 $\pm$ 0.4 1.07 $\pm$ 0.1 37.8 $\pm$ 3.2 43.5 $\pm$ 0.8 1.72 $\pm$ 0.98 43.9 $\pm$ 1.1
Walker2D-M - CQL 81.8 $\pm$ 0.7 67.7 $\pm$ 2.0 73.1 $\pm$ 0.2 77.4 $\pm$ 2.1 77.45 $\pm$ 1.66 80.1 $\pm$ 0.4
Walker2D-MR - IQL 80.2 $\pm$ 3.6 -0.2 $\pm$ 0.1 31.9 $\pm$ 1.5 78.7 $\pm$ 5.5 60.21 $\pm$ 3.8 37.5 $\pm$ 2.8
Hopper-MR - CQL 85.2 $\pm$ 8.1 0.7 $\pm$ 0.01 30.3 $\pm$ 4.7 25.1 $\pm$ 4.9 86.05 $\pm$ 13.64 19.2 $\pm$ 10.3
Sum Totals 488.4 99.27 331.4 414.3 402.6 359.4
Table 1: Performance of learning algorithms under the attacks.
To cover as diverse offline datasets as possible, here we use only one learning algorithm for each dataset. Next, we will show the case where we test the performance of different learning algorithms on the same dataset poisoned by our attack. Table 1 shows the main results of running our policy contrast attack in different learning scenarios. The naive baseline is the training with no attack. The ideal effect of the attack is the training with fully inverted rewards, as the learning algorithm tries to learn the worst-performing policy in this case. We observe that our policy contrast attack is usually effective for any combination of datasets and learning algorithms. In contrast, the random inverted attack has almost no effect on the learning result except for the Hopper-Medium Replay dataset, where both attacks are similarly effective. In many cases, our attack is also significantly more efficient than the RPI and RPP baseline attacks.
Here, we briefly discuss why the policy contrast attack should be more efficient than other baseline attacks. 1. random inverted attack: At a high level, both the policy contrast and the inverted reward attack want to inverse the agent’s ranking of good and bad policies. The inverted reward attack treats all state actions equally and inverse the performance of all policies. In contrast, the policy contrast attack focuses on the policy and the more important state actions. This could be the reason for making the policy contrast attack more efficient in terms of budget. 2. RPI and RPP attacks: The PCA attack focuses on the supported policies, while the RPI and RPP attacks focus on arbitrary policies. Since the efficient learning algorithm mostly considers only supported policies, the poisoning of such policies should have more influence on the learning process.
5.2 Universal Attack and Robustness Evaluation
To verify that our attack is universal against different learning algorithms, we take the Hopper-medium-expert dataset as an example, apply the policy contrast attack using the same setup as in the main results, and run four different learning algorithms including TD3 BC, CQL, AWAC, and ReBRAC on the corrupted dataset. The results in Figure 2 show that our attack is always effective against all four learning algorithms.
In addition, an important direct application of our attack strategy is using the attack to evaluate the robustness of an algorithm. In this example, we find that compared to other algorithms, the AWAC algorithm performs much better under our attack, suggesting that it is more robust.
Figure 2: Peformance of different learning algorithms on the same dataset under the attacks.
5.3 Influence of Attack Budget
Attack budget $B$: Here we study the influence of per-step attack budget $B$ on the attack efficiency. Denote $r_{\text{max}}=\max_{i}|r_{i}|$ as the maximal absolute reward from the dataset. In the main results, the budget required by the attacks is $B=2*r_{\text{max}}$. Here, we set the value of $B$ to be $1.5\cdot r_{\text{max}}$ and $1\cdot r_{\text{max}}$ and test on two randomly chosen datasets. The results in Fig 3 show that our policy contrast attack remains efficient with a lower budget of $B$. We note that the attack is slightly more efficient with less value of $B$. Intuitively, although the per-step corruption is less, the attack influences the performance of more policies, making it possible for the agent to learn slightly worse policies. The result also suggests that the efficiency of our attack is not significantly affected by the choice of the corruption parameters $\Delta_{1}$ and $\Delta_{2}$ at each step, as their values depend on $B$.
Figure 3: Influence of different $B$ budget on the attack.
Attack budget $C$: Here, we study the influence of attack budget $C$, the total amount of corruption, on the efficiency of the attack. We set the value of $C$ to be the amount required by a random inverted attack that corrupts $20\%$ and $40\%$ of data. The results in Fig 4 show that our attack is more efficient with a higher budget of $C$ and remains effective with a lower budget of $C$.
We also study the effect of other hyperparameters on our method. We find that the size of the good policy set has a very small influence on its efficiency. If our method only has the part of making good policy look bad or making bad policy look good, it will be much less efficient than the full method. The details of empirical results can be found in the Appendix.
Figure 4: Influence of different $C$ budget on the attack.
6 Conclusion and Limitation
In this work, we propose the first black-box reward poisoning attack in offline reinforcement learning based on our novel theoretical insights. Our empirical results suggest that the attack is effective in various learning scenarios. However, our work is limited in that we only consider black-box attacks, and the attacker requires access to the full offline training dataset.
7 Reproducibility
In the main paper, we explain the setting of the problem we study and the threat model we consider. The proofs for all theorems and lemmas can be found in the appendix. The codes we use for the experiments can be found in the supplementary materials.
Appendix A Proof for Theorems and Lemmas
Proof for Theorem 4.4 By the definition of the attack, the perturbation on the $i^{\text{th}}$ data is $\Delta_{i}=\widehat{\mathcal{R}}(s_{i},a_{i})-r_{i}$. By Assumption 4.1, the agent should learn a near-optimal policy on the reward function of the dataset, which is $\widehat{\mathcal{R}}$ under the attack. Therefore, the policy $\hat{\pi}^{*}$ learned by the agent under the attack satisfies $J_{\widehat{R}}(\hat{\pi}^{*})\geq\max_{\pi\in\Pi_{\mu}}J_{\widehat{\mathcal{R}}}(\pi)-\delta$.
Proof for Theorem B.2 The adversarial reward function constructed by the fully inverted attack is $\widehat{\mathcal{R}}=-\mathcal{R}$. By Assumption 4.1, the agent will learn a policy $\pi_{0}$ that is near-optimal on $\widehat{\mathcal{R}}$: $J_{\widehat{\mathcal{R}}}(\pi_{0})\geq\max_{\pi\in\Pi_{\mu}}J_{\widehat{\mathcal{R}}}(\pi)-\delta$. Therefore, the performance of the learned policy on the true reward function satisfies $J_{\mathcal{R}}(\pi_{0})=-J_{\widehat{\mathcal{R}}}(\pi)\leq-(\max_{\pi\in\Pi_{\mu}}J_{\widehat{\mathcal{R}}}(\pi)-\delta)=\min_{\pi\in\Pi_{\mu}}J_{\mathcal{R}}(\pi)+\delta$.
Proof for Theorem 4.7 At the last iteration of Alg 2, the performance of the policy $\pi_{K}$ learned in the iteration on the adversarial reward function ^$\widehat{\mathcal{R}}_{K}$ in the iteration is no greater than $V$ because the adversarial reward function is strictly less than the actual reward function. By Assumption 4.1, the performance of any policy $\pi\in\Pi_{\mu}$ on ^$\widehat{\mathcal{R}}_{K}$ satisfies $J_{\widehat{\mathcal{R}}_{K}}(\pi)\leq V+\delta$. For any policy $\pi\in\Pi_{\mu}:\min_{\pi_{2}\in\Pi_{2},s}||\pi(s)-\pi_{2}(s)||_{2}>d$, their performance on the actual reward function and the adversarial reward function are the same. Therefore, their actual performance must be less than $V+\delta$.
Appendix B Inverted Reward Attack
Here, we discuss the most popular offline reward poisoning attack widely considered in recent RL studies Ye et al. (2023); Zhang et al. (2022); Li et al. (2024). We start from an illustrative example first to help better understanding the attack. Following the idea of reward engineering, one can immediately find an attack that can make the agent learn the worst possible policy under Assumption 4.1.
Definition B.1 (Fully inverted reward attack).
The inverted reward attack flips the sign of all the rewards in the dataset. In Theorem B.2 we show the efficiency of the fully inverted reward attack.
Theorem B.2.
Theorem B.2 can be derived from Theorem 4.4 directly. Maximizing the cumulative reward on $-\mathcal{R}$ is equivalent to minimizing that on $\mathcal{R}$. Therefore, an efficient learning algorithm will learn the policy with the worst performance on the true reward function among the policies supported by the dataset. Despite achieving the worst outcome for the learner, the inverted reward attack faces the problem of requiring high budgets, which is impractical. Since a practical attack should only corrupt a small portion of data, we consider an attack that can only randomly invert a part of the reward. We call this attack the ‘random inverted reward attack’. This attack strategy has been widely used in previous works as a non-trivial attack Ye et al. (2023); Zhang et al. (2022).
Definition B.3 (Random inverted reward attack).
Appendix C Training Log of Main Results
Here, we show the performance of the learned policies during training in different datasets under different attacks.
Appendix D Algorithm for Learning a Set of Good Policies
@@
Appendix E Additional Ablation Study
Size of good policy set: Here, we study the influence of the size of $\Pi_{2}$ on the efficiency of the attack. We set the sizes of the good policy set to be $3,5,7,10,15$. In Fig 6, we observe that the learning outcomes of the algorithms are similar for different sizes of good policy sets. The results suggest that the number of good policies does not considerably affect our policy contrast attack.
Only Good/Bad Policies attacks: To show that it is beneficial to do both making good policies look bad and bad policies look good, we test the efficiency of the attack strategy with only one part. That is, the attack either only decreases the reward associated with the good policies or increases the reward associated with the bad policy. In Fig 7, we observe that both attacks lose some efficiency compared to the policy contrast attack. Here, we provide empirical insights on why these two attacks do not work alone. For making good policies look bad, it can happen in practice that the good policy set does not cover all policies of high performance, therefore the agent is still able to identify a high-performing policy. For making a bad policy look good, in practice, even if the bad policy already has the highest performance in the adversarial environment, without the good policy attack part, the learning agent may not be able to learn it and still converge to a good policy as the adversarial reward function is more complicated. In this case, breaking the optimality of the good policies makes it easier for the learning agent to converge to the bad policy.